• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name markowitz

    Log entry time 17:48:47 on January17,2004

    Entry number 116013

    keyword=shift summary

    began shift with end of carbon running (run 1866).

    Performed septum cooling from 08:00 to 11:30

    Ran 12Carbon e,e'K, varying current for hall c

    Run 1867 50 uA 500K events
    run 1868 75 uA 500K events RICH tripped at beginning of run; reset
    run 1869 100 uA 500K events
    run 1870 100 uA 500K events
    run 1871 100 uA 500K events
    run 1872 100 uA 500K events
    run 1873 8 uA 175K events harp scans/SLI work beginning. MCC is playing with beam.

    15:30 MCC is done with beam studies of SLI/harp scans. Ready to restore beam to 100 uA.

    run 1874 100 uA